# Scrutinize What We Ignore: Reining In Task Representation Shift Of Context-Based Offline Meta Reinforcement Learning

### v2

@@ Abstract
Offline meta reinforcement learning (OMRL) has emerged as a promising approach for interaction avoidance and strong generalization performance by leveraging pre-collected data and meta-learning techniques. Previous context-based approaches predominantly rely on the intuition that alternating optimization between the context encoder and the policy can lead to performance improvements, as long as the context encoder follows the principle of maximizing the mutual information between the task and the task representation ($I(Z;M)$) while the policy adopts the standard offline reinforcement learning (RL) algorithms conditioning on the learned task representation. Despite promising results, the theoretical justification of performance improvements for such intuition remains underexplored. Inspired by the return discrepancy scheme in the model-based RL field, we find that the previous optimization framework can be linked with the general RL objective of maximizing the expected return, thereby providing a feasible explanation concerning performance improvements. Furthermore, after scrutinizing this optimization framework, we find it ignores the impacts stemming from the variation of the task representation in the alternating optimization process, which may lead to performance improvement collapse. We name this issue task representation shift and theoretically prove that the monotonic performance improvements can be guaranteed with appropriate context encoder updates. We set different manners to rein in the task representation shift on three widely adopted training objectives concerning maximizing $I(Z;M)$ across different data qualities. Empirical results show that reining in the task representation shift can indeed improve performance. Our work opens up a new avenue for OMRL, leading to a better understanding between the performance and the task representation.
@@ 1 Introduction
Figure 1: Our training framework compared to the previous training framework. They both adopt the alternating optimization framework to train the context encoder and the policy. However, ours considers the variation of task representation ignored by the previous training framework.
Intriguingly, we formalize this training framework as an alternating two-stage optimization framework and then link it with the general RL objective of maximizing the expected return. To be specific, maximizing $I(Z;M)$ and adopting standard offline RL algorithms can be interpreted as consistently raising the lower bound of the expected return conditioning on the optimal task representation distribution among the expectation of tasks. This is achieved by extending the return discrepancy scheme [janner2019trust] to the COMRL framework. Thus, this analysis provides a feasible explanation for the performance improvement guarantee.
More importantly, after scrutinizing this optimization framework, we find it ignores the impacts stemming from the variation of the task representation in the alternating process. Variations of the task representation that are either too large or too small can lead to performance improvement collapse. To highlight the characteristic, we name this issue task representation shift and theoretically prove that it is possible to achieve monotonic performance improvement guarantees with proper updates.
To show the impacts of this issue, we set different manners to rein in the task representation shift on contrastive objective adopted in [li2020focal, gao2024context, li2024towards] and reconstruction objective adopted in [zintgraf2019varibad, dorfman2021offline, li2024towards] across different data qualities. These two objectives are proven to maximize the approximate bounds of $I(Z;M)$ [li2024towards]. We also introduce the cross-entropy objective, which is a straightforward discrete approximation to $I(Z;M)$. Empirical results show that reining in the task representation shift can indeed improve performance. Our work opens up a new avenue for OMRL, leading to a better understanding between the performance and the task representation.
@@ 2 Related Works
Context-based Offline Meta RL. As the marriage between context-based meta RL and offline RL, COMRL combines the merits of both sides. Specifically, COMRL methods [li2020focal, yuan2022robust, gao2024context, li2024towards] leverage the offline dataset to train a context encoder to learn robust task representations, and then pass the representation to the policy and value function as input. At test time, COMRL methods leverage the generalization ability of the learned context encoder to perform meta-adaptation. Acknowledging the theoretical insights from [li2024towards], prior COMRL works [li2020focal, yuan2022robust, gao2024context, li2024towards] mainly focus on how to design a context encoder learning algorithm to better approximate $I(Z;M)$. However, our work moves the focus on how to get better asymptotic performance based on a given context encoder learning algorithm. Center around this motivation, we identify a new issue called task representation shift, which is ignored by the previous COMRL endeavors.
@@ 3.2 Context-based OMRL
Previous COMRL endeavors hold that the variation over different tasks can be represented by a compact representation $z$ [yuan2022robust]. In practice, they choose to train a context encoder $Z(\cdot|x)$ to extract the task information [li2020focal, dorfman2021offline], where $x$ denotes the given context. Then, the policy is learned by conditioning on the task representation as $\pi(a|s,Z(\cdot|x);\theta)$. Naturally, there exists an optimal task representation distribution $Z^{*}(\cdot|x)$ that meets the optimal return for any policy parameter $\theta$ is determined by $\pi(a|s,Z^{*}(\cdot|x);\theta)$. Hence, the objective of COMRL is formalized as:
@@
Previous COMRL works [li2020focal, yuan2022robust, gao2024context, li2024towards] are proven to optimize the context encoder by maximizing the approximate bounds of $I(Z;M)$, as shown in Theorem 3.1. They hold the intuition that using such kind of manner to optimize the context encoder and adopting the standard offline RL algorithms to optimize the policy can lead to performance improvements. However, the theoretical justification of performance improvement for such intuition has been less explored.
@@ 4.1 A Performance Improvement Perspective Towards Prior Works
Our goal is to induce an optimization objective served for COMRL to get the performance improvement guarantee. Motivated by the return discrepancy scheme [janner2019trust] in the model-based RL field, we can similarly build a tractable lower bound for $J^{*}(\theta)$.
@@
From Definition 4.2, we can know that if the estimation error $|J^{*}(\theta)-J(\theta)|$ can be upper-bounded, the lower bound for $J^{*}(\theta)$ can be established simultaneously. Then, Eq. (4) can be interpreted as a unified training objective for both the context encoder and the policy. Specifically, we can alternate the optimization target to lift the lower bound for $J^{*}(\theta)$, where the holistic optimization process can be seen as a two-stage alternating framework with the first stage updating the context encoder to minimize $|J^{*}(\theta)-J(\theta)|$ and the second stage updating the policy to maximize $J(\theta)-|J^{*}(\theta)-J(\theta)|$. Hence, the performance improvement guarantee can be achieved.
@@
Denote $Z_{mutual}(\cdot|x)$ as the optimal task representation distribution served for maximizing $I(Z;M)$  [^1] Whether the solution of maximizing $I(Z;M)$ is equivalent to the solution of maximizing the expected return is still an open problem. Hence, we introduce an extra notation here.. Assume the reward function is upper-bounded by $R_{max}$. For an arbitrary policy parameter $\theta$, when meeting Assumption 4.1, the return bound in COMRL can be formalized as:
@@
Theorem 4.3 and Eq. (6) indicate the following two principles for the context encoder and the policy. For the context encoder learning, we should minimize the return bound, namely minimizing $\mathbb{E}_{m,x}|Z(\cdot|x)-Z_{mutual}(\cdot|x)|$. For the policy learning, we should lift the lower bound of $J^{*}(\theta)$, namely maximizing $J(\theta)-\frac{2R_{max}L_{z}}{(1-\gamma)^{2}}\mathbb{E}_{m,x}(|Z(\cdot|x)-Z_{mutual}(\cdot|x)|+|Z_{mutual}(\cdot|x)-Z^{*}(\cdot|x)|)\equiv J(\theta)$ as the task representation related terms are constants for policy optimization. To maximize $J(\theta)$, we can adopt standard offline RL algorithms [wu2019behavior, fujimoto2021minimalist, kumar2020conservative] similar to [yang2022unified].
Recall that prior COMRL works adopt alternatively training the context encoder by maximizing $I(Z;M)$ and training the policy network by conditioning on the learned task representation as well as applying the standard offline RL algorithms. As maximizing $I(Z;M)$ can be approximately seen as minimizing $\mathbb{E}_{m,x}|Z(\cdot|x)-Z_{mutual}(\cdot|x)|$, we argue that prior COMRL works can be interpreted by this optimization framework, thus filling the gap between the intuition and the theoretical justification.
However, this optimization framework only considers the discrepancy between $Z(\cdot|x)$ and $Z_{mutual}(\cdot|x)$, leaving the variation of the task representation alone. In the next section, we will show that this may lead to performance improvement collapse.
@@ 4.2 Monotonic Performance Improvement Concerning Task Representation Shift
We attempt to build a theoretical framework that can achieve monotonic performance improvement. Hence, we turn to use the framework of performance difference bound.
@@
According to Definition 4.4, we need to find the positive $C$ to improve the performance monotonically. We denote $J^{mutual}(\theta_{i}),i\in[1,2]$ as the expected return of the policy $\pi(\cdot|s,Z_{mutual}(\cdot|x);\theta_{i})$ among the expectation of tasks and then introduce a mild assumption for the task representation distribution.
Assumption 4.5.
The support set of $Z_{i}(\cdot|x),i\in[1,2]$ covers the support set of $Z_{mutual}(\cdot|x)$.
Assumption 4.5 usually holds in practice as when we give $z\sim Z_{mutual}(\cdot|x)$ to the task representation distribution $Z_{i}(\cdot|x)$ parameterized by the neural network, we can always get a non-zero probability. Given that augmented state $s^{+}=[s,z]$, by applying standard offline RL algorithms with enough samples, the policy update would lead to performance improvement when meeting Assumption 4.5, which means
$\epsilon^{mutual}_{12}\triangleq J^{mutual}(\theta_{2})-J^{mutual}(\theta_{1})>0$ (8)
By leveraging Definition 4.4 and the above assumptions, we derive the lower bound of the performance difference.
@@
Assume the reward function is upper-bounded by $R_{max}$. When meeting Assumption 4.1 and Assumption 4.5, the lower bound of the performance difference in COMRL can be formalized as:
$J^{2}(\theta_{2})-J^{1}(\theta_{1})\geq\epsilon^{mutual}_{12}-\frac{2R_{max}L_{z}}{(1-\gamma)^{2}}\mathbb{E}_{m,x}[2|Z_{2}(\cdot|x)-Z_{mutual}(\cdot|x)|+|Z_{2}(\cdot|x)-Z_{1}(\cdot|x)|]$ (9)
@@
$|Z_{2}(\cdot|x)-Z_{mutual}(\cdot|x)|\leq\frac{(1-\gamma)^{2}}{4R_{max}L_{z}}\epsilon^{mutual}_{12}-\frac{1}{2}|Z_{2}(\cdot|x)-Z_{1}(\cdot|x)|$ (10)
The above equation unveils the key influence of the variation of task representation $|Z_{2}(\cdot|x)-Z_{1}(\cdot|x)|$ for two aspects. If $|Z_{2}(\cdot|x)-Z_{1}(\cdot|x)|$ is too small, $|Z_{2}(\cdot|x)-Z_{mutual}(\cdot|x)|$ may greater than the right-hand side of Eq. (10). On the contrary, if $|Z_{2}(\cdot|x)-Z_{1}(\cdot|x)|$ is too large, the right-hand side of Eq.(10) may be negative. Both of these cases would lead to the performance improvement collapse. Thus, we argue that the impacts of $|Z_{2}(\cdot|x)-Z_{1}(\cdot|x)|$ cannot be ignored.
To highlight the impacts of $|Z_{2}(\cdot|x)-Z_{1}(\cdot|x)|$, we name this issue task representation shift. Under some mild assumptions, we can conclude that it is possible to achieve monotonic performance improvement on the condition of proper context encoder updates, as shown in Theorem 4.10.
@@
The impacts of task representation shift can be upper-bounded by $\beta$.
@@
There exists an $\alpha$ for any given $Z_{1}(\cdot|x)$, $|Z_{2}(\cdot|x)-Z^{empirical}_{2}(\cdot|x)|^{2}\leq\frac{\alpha}{b}$, where $Z_{2}(\cdot|x)$ denotes the context encoder updated by maximizing $I(Z;M)$ with the data size $b$ from the training dataset based on $Z_{1}(\cdot|x)$ and $Z^{empirical}_{2}(\cdot|x)$ denotes the context encoder updated by fitting the empirical distribution on $b$ i.i.d samples from $Z_{mutual}(\cdot|x)$ based on $Z_{1}(\cdot|x)$  [^2] We set discrepancy to be subject to an inverse proportional function is because as the training samples increase, $Z_{2}$ and $Z^{empirical}_{2}$ should become similar..
@@
Denote $\kappa$ as $\frac{(1-\gamma)^{2}}{4R_{max}L_{z}}\epsilon^{mutual}_{12}-\frac{1}{2}\beta$ and $vol(\mathbb{Z})$ as the volume of the task representation coverage simplex. Given that the context encoder has already been trained by maximizing $I(Z;M)$ to some extent. When meeting Assumption 4.1, Assumption 4.5, Assumption 4.7, Assumption 4.8 and Assumption 4.9, with a probability greater than $1-\xi$, we can get the monotonic performance improvement guarantee by updating the context encoder with maximizing $I(Z;M)$ from at least extra $k$ samples, where:
@@
Theorem 4.10 shows the connection between the needed update data size $k$ and the performance improvement brought only by the policy update $\epsilon^{mutual}_{12}$. If the calculated $k$ is larger than the given batch size, the context encoder should avoid this update and wait for the accumulation of $\epsilon^{mutual}_{12}$. As $\epsilon^{mutual}_{12}$ increases, $k$ decreases. When $k$ is smaller than the given batch size, the monotonic performance improvement can be achieved by updating the context encoder. Notice that this updating process can be performed multiple times, as long as satisfying Assumption 4.7. The updating process is visualized in Figure 1 right. Compared to the original alternating framework, Theorem 4.10 unveils that the core part for better performance is to adjust the update of the context encoder to rein in the task representation shift, thereby showcasing the advantage over the previous works. The general algorithmic framework is shown in Algorithm 1, where the way to adjust the update of the context encoder is colored by red.

### v3

@@ Abstract
Offline meta reinforcement learning (OMRL) has emerged as a promising approach for interaction avoidance and strong generalization performance by leveraging pre-collected data and meta-learning techniques. Previous context-based approaches predominantly rely on the intuition that alternating optimization between the context encoder and the policy can lead to performance improvements, as long as the context encoder follows the principle of maximizing the mutual information between the task variable $M$ and its latent representation $Z$ ($I(Z;M)$) while the policy adopts the standard offline reinforcement learning (RL) algorithms conditioning on the learned task representation. Despite promising results, the theoretical justification of performance improvements for such intuition remains underexplored. Inspired by the return discrepancy scheme in the model-based RL field, we find that the previous optimization framework can be linked with the general RL objective of maximizing the expected return, thereby explaining performance improvements. Furthermore, after scrutinizing this optimization framework, we find it ignores the variation of the task representation in the alternating optimization process, which weakens the condition necessary for monotonic performance improvements, and may therefore violate the monotonicity. We name this issue task representation shift and theoretically prove that the monotonic performance improvements can be guaranteed with appropriate context encoder updates. We use different settings to rein in the task representation shift on three widely adopted training objectives concerning maximizing $I(Z;M)$ across different data qualities. Empirical results show that reining in the task representation shift can indeed improve performance. Our work opens up a new avenue for OMRL, leading to a better understanding between the task representation and performance improvements.
@@ 1 Introduction
Figure 1: Our training framework compared to the previous training framework. They both adopt the alternating optimization framework to train the context encoder and the policy. However, ours considers the variation of task representation ignored by the previous training framework by imposing the extra condition to determine whether needs to update the context encoder.
Intriguingly, we formalize this training framework as an alternating two-stage optimization framework and then link it with the general RL objective of maximizing the expected return. To be specific, maximizing $I(Z;M)$ and adopting standard offline RL algorithms can be interpreted as consistently raising the lower bound of the expected return conditioning on the optimal task representation distribution among the expectation of tasks (See Section 4.1). This is achieved by extending the return discrepancy scheme [janner2019trust] to the COMRL framework. Thus, this analysis provides a feasible explanation for the performance improvement guarantee.
More importantly, after scrutinizing this optimization framework, we find it ignores the impacts stemming from the variation of the task representation in the alternating process. This would cause the optimization framework to incorrectly conclude that the monotonic performance improvement can be guaranteed with a better approximation to the optimal task representation distribution. By explicitly modeling the variation of task representation, we prove that it is a critical part of monotonic performance improvement. Without this, the monotonicity can be violated. To highlight the characteristic, we name this issue task representation shift and theoretically prove that it is possible to achieve monotonic performance improvement with appropriate updates (See Section 4.2).
To show the impacts of this issue, we use different settings to rein in the task representation shift on contrastive objective adopted in [li2020focal, gao2024context, li2024towards] and reconstruction objective adopted in [zintgraf2019varibad, dorfman2021offline, li2024towards] across different data qualities. These two objectives are proven to maximize the approximate bounds of $I(Z;M)$ [li2024towards]. We also introduce the cross-entropy objective, which is a straightforward discrete approximation to $I(Z;M)$. Empirical results show that reining in the task representation shift can indeed improve performance. Our work opens up a new avenue for OMRL, leading to a better understanding between the task representation and performance improvements.
@@ 2 Related Works
Context-based Offline Meta RL. As the marriage between context-based meta RL and offline RL, COMRL combines the merits of both sides. Specifically, COMRL methods [li2020focal, yuan2022robust, gao2024context, li2024towards] leverage the offline dataset to train a context encoder to learn robust task representations, and then pass the representation to the policy and value function as input. At test time, COMRL methods leverage the generalization ability of the learned context encoder to perform meta-adaptation. Acknowledging the theoretical insights from [li2024towards], prior COMRL works [li2020focal, yuan2022robust, gao2024context, li2024towards] mainly focus on how to design a context encoder learning algorithm to better approximate $I(Z;M)$. However, our work moves the focus on how to refine the condition of monotonic performance improvements based on a given context encoder learning algorithm. Center around this motivation, we identify a new issue called task representation shift, which is ignored by the previous COMRL endeavors.
@@ 3.2 Context-based OMRL
Previous COMRL endeavors hold that the variation over different tasks can be represented by a compact representation $z$ [yuan2022robust]. In practice, they choose to train a context encoder $Z(\cdot|x)$ to extract the task information [li2020focal, dorfman2021offline], where $x$ denotes the given context. Then, the policy is learned by conditioning on the task representation as $\pi(a|s,Z(\cdot|x);\theta)$. Naturally, there exists an optimal task representation distribution $Z^{*}(\cdot|x)$ that meets the optimal expected return for any policy parameter $\theta$ is determined by $\pi(a|s,Z^{*}(\cdot|x);\theta)$. Hence, the objective of COMRL is formalized as:
@@
Previous COMRL works [li2020focal, yuan2022robust, gao2024context, li2024towards] are proven to optimize the context encoder by maximizing the approximate bounds of $I(Z;M)$, as shown in Theorem 3.1. They hold the intuition that using such kind of approach to optimize the context encoder and adopting the standard offline RL algorithms to optimize the policy can lead to performance improvements. However, the theoretical justification of performance improvement for such intuition has been less explored. For later use, we denote $Z_{mutual}(\cdot|x)$ as the optimal solution for maximizing $I(Z;M)$. Whether $Z_{mutual}(\cdot|x)$ is equivalent to $Z^{*}(\cdot|x)$ is still an open problem. Hence, we introduce an extra notation here to avoid confusion.
@@ 4.1 A Performance Improvement Perspective Towards Prior Works
Our goal is to arrive at an optimization objective for COMRL to get the performance improvement guarantee. Motivated by the return discrepancy scheme [janner2019trust] in the model-based RL field, we can similarly build a tractable lower bound for $J^{*}(\theta)$.
@@
From Definition 4.2, we know that if the estimation error $|J^{*}(\theta)-J(\theta)|$ can be upper-bounded, the lower bound for $J^{*}(\theta)$ can be established simultaneously. Then, Eq. (4) can be interpreted as a unified training objective for both the context encoder and the policy. Specifically, we can alternate the optimization target to lift the lower bound for $J^{*}(\theta)$, where the holistic optimization process can be seen as a two-stage alternating framework with the first stage updating the context encoder to minimize $|J^{*}(\theta)-J(\theta)|$ and the second stage updating the policy to maximize $J(\theta)-|J^{*}(\theta)-J(\theta)|$ as $J^{*}(\theta)\geq J(\theta)-|J^{*}(\theta)-J(\theta)|$. Hence, the performance improvement guarantee can be achieved.
@@
Assume the reward function is upper-bounded by $R_{max}$. For an arbitrary policy parameter $\theta$, when meeting Assumption 4.1, the return bound in COMRL can be formalized as:
@@
Theorem 4.3 and Eq. (6) indicate the following two principles for the context encoder and the policy. For the context encoder learning, we should minimize the return bound, namely minimizing $\mathbb{E}_{m,x}|Z(\cdot|x)-Z_{mutual}(\cdot|x)|$. For the policy learning, we should lift the lower bound of $J^{*}(\theta)$, namely maximizing $J(\theta)-\frac{2R_{max}L_{z}}{(1-\gamma)^{2}}\mathbb{E}_{m,x}(|Z(\cdot|x)-Z_{mutual}(\cdot|x)|+|Z_{mutual}(\cdot|x)-Z^{*}(\cdot|x)|)\equiv J(\theta)$, as the task representation related terms are constants for policy optimization. To maximize $J(\theta)$, we can adopt standard offline RL algorithms [wu2019behavior, fujimoto2021minimalist, kumar2020conservative] similar to [yang2022unified].
Recall that previous COMRL works adopt alternatively training the context encoder by maximizing $I(Z;M)$, which can be approximately seen as minimizing $\mathbb{E}_{m,x}|Z(\cdot|x)-Z_{mutual}(\cdot|x)|$ and training the policy by conditioning on the learned task representation as well as applying the standard offline RL algorithms. Therefore, we argue that previous COMRL works can be interpreted by this framework, thus providing an explanation concerning performance improvement guarantee.
However, this optimization framework only considers the discrepancy between $Z(\cdot|x)$ and $Z_{mutual}(\cdot|x)$, without considering the variation of $Z(\cdot|x)$. In the next section, we will show that this may violate the monotonicity of the performance improvements.
@@ 4.2 Monotonic Performance Improvement Concerning Task Representation Shift
In this section, we aim to demonstrate that as the previous optimization framework in Section 4.1 doesn’t model the variation of task representation explicitly, this framework would provide an insufficient condition for monotonic performance improvement. We first show conditions for monotonic performance improvement of this framework.
Corollary 4.4 (Monotonic performance improvement condition for previous COMRL works).
When meeting Assumption 4.1, the condition for monotonic performance improvement of previous COMRL works is:
$\epsilon^{*}_{12}\triangleq J^{*}(\theta_{2})-J^{*}(\theta_{1})\geq\frac{4R_{max}L_{z}}{(1-\gamma)^{2}}\mathbb{E}_{m,x}(|Z(\cdot|x)-Z^{*}(\cdot|x)|)$ (7)
Proof: See Section 8.2 in the Appendix. ∎
As shown in Corollary 4.4, the monotonic performance improvement can be guaranteed with only better approximation to $Z^{*}(\cdot|x)$. Next, we will introduce the performance difference bound framework to model variation of task representation.
@@
According to Definition 4.5, we need to find the positive $C$ to improve the performance monotonically. To achieve this, we can derive the lower bound of the performance difference.
@@
Assume the reward function is upper-bounded by $R_{max}$. When meeting Assumption 4.1, the lower bound of the performance difference in COMRL can be formalized as:
$J^{2}(\theta_{2})-J^{1}(\theta_{1})\geq\epsilon^{*}_{12}-\frac{2R_{max}L_{z}}{(1-\gamma)^{2}}\mathbb{E}_{m,x}[2|Z_{2}(\cdot|x)-Z^{*}(\cdot|x)|+|Z_{2}(\cdot|x)-Z_{1}(\cdot|x)|]$ (9)
@@
$\epsilon^{*}_{12}-\frac{2R_{max}L_{z}}{(1-\gamma)^{2}}\mathbb{E}_{m,x}[2|Z_{2}(\cdot|x)-Z^{*}(\cdot|x)|+|Z_{2}(\cdot|x)-Z_{1}(\cdot|x)|]\geq 0$ (10)
Compared to the condition in Corollary 4.4, Eq. 10 presents an additional term $|Z_{2}(\cdot|x)-Z_{1}(\cdot|x)|$ to achieve monotonic performance improvement, which corresponds exactly to the previous ignored impacts stemming from the variation of the task representation. To achieve monotonic performance improvements, the optimization of task representation should consider not only the approximation to $Z^{*}(\cdot|x)$, but also the magnitude of the update. For example, if we assume that $Z_{1}(\cdot|x)$ is trained from scratch and the update process from $Z_{1}(\cdot|x)$ to $Z_{2}(\cdot|x)$ brings $Z_{2}(\cdot|x)$ close to $Z^{*}(\cdot|x)$, then with the condition in Corollary 4.4, the monotonic performance improvement can be easily achieved with small $\epsilon^{*}_{12}$. However, for the condition in Eq. (10), small $\epsilon^{*}_{12}$ may cause the violation of monotonicity as $|Z_{1}(\cdot|x)-Z_{2}(\cdot|x)|$ remains large.
To highlight the impacts of $|Z_{2}(\cdot|x)-Z_{1}(\cdot|x)|$, we name this issue Task Representation Shift. Under some mild assumptions, we can conclude that it is possible to achieve monotonic performance improvement with sufficient policy improvement $\epsilon^{*}_{12}$ on the condition of appropriate encoder updates, as shown in Theorem 4.10.
@@
The impacts of task representation shift are upper-bounded by $\beta$ and less than the policy improvement $\epsilon^{*}_{12}$ with a certain coefficient.
@@
There exists an $\alpha$ for any given $Z_{1}(\cdot|x)$, $|Z_{2}(\cdot|x)-\tilde{Z}_{2}(\cdot|x)|^{2}\leq\frac{\alpha}{b}$, where $Z_{2}(\cdot|x)$ denotes the context encoder updated by maximizing $I(Z;M)$ with the data size $b$ randomly sampled from the training dataset based on $Z_{1}(\cdot|x)$ and $\tilde{Z}_{2}(\cdot|x)$ denotes the context encoder updated by fitting the empirical distribution on $b$ i.i.d samples from $Z^{*}(\cdot|x)$ based on $Z_{1}(\cdot|x)$  [^1] We set discrepancy to be subject to an inverse proportional function is because as the training samples increase, $Z_{2}$ and $\tilde{Z}_{2}$ should become similar..
@@
Denote $\kappa$ as $\frac{(1-\gamma)^{2}}{4R_{max}L_{z}}\epsilon^{*}_{12}-\frac{1}{2}\beta$ and $vol(\mathbb{Z})$ as the volume of the task representation coverage simplex. Given that the context encoder has already been trained by maximizing $I(Z;M)$ to some extent. When meeting Assumption 4.1, 4.7, 4.8 and 4.9, with a probability greater than $1-\xi$, we can get the monotonic performance improvement guarantee by updating the context encoder via maximizing $I(Z;M)$ from at least extra $k$ samples, where:
@@
Theorem 4.10 shows the connection between the needed update data size $k$ and the performance improvement brought only by the policy update $\epsilon^{*}_{12}$. If the calculated $k$ is larger than the given batch size, the context encoder should avoid this update and wait for the accumulation of $\epsilon^{*}_{12}$. As $\epsilon^{*}_{12}$ increases, $k$ decreases. When $k$ is smaller than the given batch size, the monotonic performance improvement can be achieved by updating the context encoder. Notice that this updating process can be performed multiple times, as long as satisfying Assumption 4.7. The updating process is visualized in Figure 1 right. Compared to the original alternating framework, Theorem 4.10 unveils that the core part for better performance improvement is to adjust the update of the context encoder to rein in the task representation shift, thereby showcasing the advantage over the previous works. The general algorithmic framework is shown in Algorithm 1, where the way to adjust the update of the context encoder is colored by red.